13. A project has an initial cost of $38,300 and anticipated cash flows of $9,200, $18,700,...

13.

A project has an initial cost of $38,300 and anticipated cash flows of $9,200, $18,700, $14,600 for Years 1 to 3, respectively. What is the profitability index value if the required return is 9.5 percent?

.99

.86

.92

1.16

1.09
Answer #1

Calculation of profitability index:

Profitability index= present value of cash inflows/initial outlay

Present value of cash inflows= (9200*pvf@9.5% for 1 year)+ (18700*pvf@9.5% for 2 year)+(14600*pvf@9.5% for year3)

Present value of cash inflows= (9200*0.913)+(18700*0.834)+(14600*0.762)

Present value of cash inflows= 8399.6+15595.8+11125.2= 35120.6

Profitability index= 35120.6/38300= 0.92

So correct answer is .92